Can you please solve this equation:
2sqrt(x + 8) = 3sqrt(x - 2)

Answer by richwmiller(17219)

2sqrt(x + 8) = 3sqrt(x - 2)
divide both sides by 2
sqrt(x + 8) = 3sqrt(x - 2)/2
square both sides
x+8=9(x-2)/4
expand terms on right
x+8=9x/4-9/2
subtract 9x/4-9/2 from both sides
4/x+32/4-9x/4+18/4=0
combine
50/4-5x/4=0
add 5x/4 to both sides
50/4=5x/4
multiply by 4
50=5x
10=x
check
2sqrt(x + 8) = 3sqrt(x - 2)
2sqrt(18)=3sqrt(8)
2sqrt(9*2)=3sqrt(4*2)
3*2sqrt(2)=2*3sqrt(2)
ok